How do you solve x/5 + 3= 7
Answer by mszlmb(115)

thnx 4 correcting me; just comes 2 show, u really get this stuff after all!
x/5+3= 7
-3 on both sides
x/5+3 -3=7 -3
simplify
x/5=4
multiply by 5 on both sides
x/5 *5=4 *5
simplify
x=20